if 100cm of oxygen diffusion in 4 seconds and 500cm3 of gas diffused in 3 seconds.calculate the rmm(o=10)
1
Expert's answer
2018-02-23T07:05:38-0500
The relative molar mass (rmm) can be calculated from the Graham’s law of diffusion:
(Rate A / Rate B )2 = rmm B / rmm A,
where the rate of diffusion the volume diffused per second:
Rate Oxygen = 100 cm3 / 4 s = 25 cm3/s,
Rate Gas = 500 cm3 / 3 s = 166.7 cm3/s.
Since the molar mass of oxygen is 32 g/mol, then the molar mass of a gas is:
rmm Gas = 32 g/mol (25/166.7)2 = 0.72 g/mol.
This answer is not reasonable, as the lightest gas has a molar mass of 2 g/mol (hydrogen). Perhaps, there is a mistake in the given values.
Answer: 0.72 g/mol.